l. Armbrister Windows makes windows which have a production standard of an average of 0.375 inches thick with a standard deviation of 0.050 inches. A random sample of 64 windows yields an average of 0.382 inches. Calculate the probability of obtaining a random sample of 64 windows with an average of at least 0.382 inches. 2. The time it takes a mechanic to replace a brake pad is known to be normally distributed with a mean of 28 minutes and a standard deviation of 10 minutes. What size sample should be taken if the goal is to get the standard deviation of the sampling mean to be 3 minutes? 3. It is estimated that 82%ofall computer uses in Edina, Mn, make at least one purchase p month through the internet. In a random sample of 100 computer users from Edina, Mn, 74o f them made an internet purchase in the past month. Find the probability of finding no more than 74 computer users from Edina, MN making an internet purchase from a sample of 100. 4. Marketing research indicates that 37% of all customers at a nationwide pizza chain are college students. A random sample of 50 customers at a local franchise yields 15 college student customers. Find the probability that a random sample of 50 customers would yield between 14 and 20 college student customers.
please let me know and way to answer
Answer #1

Solution (1):

     We are given that : Armbrister windows makes windows which have a production standard of an average of 0.375 inches thick with a standard deviation of 0.050 inches.

Thus Mean = \mu = 0.375 and standard deviation = \sigma = 0.050

A random sample of 64 windows yields an average of 0.382 inches.
We have to find the probability that : a random sample of 64 windows with an average of at least 0.382 inches.

That is we have to find : P(\bar{x}\geq 0.382)=.......?

Since sample size n = 64 > 30 , we can assume large sample and hence using Central limit theorem , sampling distribution of sample mean has an approximate Normal distribution with mean of sample means = \mu_{\bar{x}} = \mu = 0.375 and standard deviation of sample means = \sigma_{\bar{x}} = \sigma / \sqrt{n} = 0.050/ \sqrt{64} = 0.050 / 8 = 0.00625

Thus to find : P(\bar{x}\geq 0.382)=.......? find z score .

z= \frac{\bar{x}-\mu_{\bar{x}}}{\sigma_{\bar{x}}}

z= \frac{0.382-0.375}{0.00625}

z= \frac{0.007 }{0.00625}

z= 1.12

Thus we get :

P(\bar{x}\geq 0.382)=P(z\geq 1.12)

P(\bar{x}\geq 0.382)=1-P(z\leq 1.12)

Look in z table for z = 1.1 and 0.02 and find corresponding area.

100 oil 02 .05 .06 07 08 .09 0.0 .5000 .5040 . 0.1 .5398 .5438 0.2 .5793 5 5080 .5120 5160 5199 5239 5279 5319 5359 5517 5557

P( Z < 1.12 ) = 0.8686

Thus

P(\bar{x}\geq 0.382)=1-P(z\leq 1.12)

P(\bar{x}\geq 0.382)=1-0.8686

P(\bar{x}\geq 0.382)=0.1314

Thus the probability that a random sample of 64 windows with an average of at least 0.382 inches is 0.1314 .